ou invested $7,000 in a savings deposit 6 quarters ago and it has grown to $7887 today. What nominal rate of annual interest (compounding quarterly) did you earn?
Answer #1

Interest is compounded quarterly.

Let quarterly interest rate be x

Then, 7,000(1+x)6 = 7,887

(1+x)6 = 1.1267

(1+x) = (1.1267)1/6

1+x = 1.020083

X = 0.020083

Annual Interest = 0.020083*4 = 8.03% (approx)